# Online Flow Time Minimization: Tight Bounds for Non-Preemptive Algorithms

### v2

@@ 8.3 Generalization to Multiple Machines with Kill-and-Restart
Any randomized online algorithm for minimizing the total flow time on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed pair $(n,m)$, where $\Gamma(n,m)=\Omega(\sqrt{n/m})$, even with the capability of kill-and-restart.
@@
By Lemma 23, $\textup{{OPT}}=O(km)$, hence any deterministic algorithm is $\Omega(k)=\Omega(n/m)$ competitive and the theorem is proved. ∎
@@ 8.4 Deterministic Algorithms: Single Machine with Kill-and-Restart
    if ALG hasn’t kill job $u$ before time $t^{\prime}+\tau$ then
@@
All deterministic algorithms with kill-and-restart on single-machine total flow time minimization have a competitive ratio of at least $\Omega(n/\log n)$.
@@ 8.5 Deterministic Algorithms: Multiple Machines without Kill-and-Restart
The $\Omega(n/m^{2}+\sqrt{n/m})$ follows by combining our randomized lower bound of $\Omega(\sqrt{n/m})$ with the deterministic $\Omega(n/m^{2})$ lower bound from DBLP:journals/dam/EpsteinS06, as presented below.
@@
Any deterministic online algorithm for minimizing the total flow time on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed $n$ and $m$, where $\Gamma(n,m)=\Omega(n/m^{2}+\sqrt{n/m})$.
@@
We first show that for every $n$, $m$, we can construct an adversary to prove $\Omega(n/m^{2})$, the same as DBLP:journals/dam/EpsteinS06. We still present it for completeness. Consider the following case for any deterministic algorithm ALG:
@@
OPT can start all small jobs at their release time immediately; it can also solve the first job at 0 if $t>1$, or solve the first job at $2$ otherwise. Hence, OPT is $O(1)$ in both cases. For ALG, as it can only use m1m-1 machines to solve the small jobs, at most m1m-1 jobs can be solved between the releases of two small job batchs. The total flow time for ALG is at least $\sum_{i=1}^{n/m}i/n=\Omega(n/m^{2})$. Therefore, the competitive ratio is $\Omega(n/m^{2})$.
@@ 9 Lower Bounds for Algorithms without the Knowledge of $n$
Adversary construction. We construct an oblivious adversary as follows. At time 0, the adversary releases a single job of size $2$. For any algorithm, let $p(t)$ denote the probability density function that ALG starts this job at time $t$ if only the single job is present, which is known by the oblivious adversary since the probability distribution solely depend on the description of the algorithm and does not depend on algorithm’s randomized results. The algorithm must fall into one of the following categories (the parameters $x$ and $y$ will be fixed later):
@@
    For all $t\in\{1,2,\dots,x\}$, it holds that $\int_{t}^{t+1}p(t)\,dt<y$. In this case, the probability that the size-$2$ job has not been started by time $x$ is at least 1xy1-xy. The adversary releases $n_{0}$ dummy jobs of size 0 at time $x+1$. If ALG delays the size-$2$ job until after time $x$, each dummy job has flow time at least $x$, while $\textup{{OPT}}\leq 2$. Therefore,
@@
It is known that even when $n$ is known in advance, there exists a lower bound of $\Omega(\sqrt{n/m})$. Therefore, it suffices to generalize Theorem 9 to obtain a hard instance with ratio $\Omega(n/m^{2})$, from which the theorem follows.
Adversary construction. We use the same setup as in the single-machine case: the adversary first releases a size-2 job at time 0 and observes the algorithm’s random behavior. The only difference lies in how we construct the hard instance for Type A algorithms in the multi-machine setting.
Hard instance for Type A.
If ALG is of Type A, the adversary releases $n_{1}$ additional jobs of size $1/k$, where $k=n_{1}/m$. The release pattern is as follows:
@@
Appendix A Schedule Small Jobs: Non-Preemptive Shortest Job First
@@
Next, for any threshold $p$ and time $t$, we define $J_{\leq p}(t)$ as the set of jobs with $p_{j}\leq p$ and release time $r_{j}\leq t$. We are interested in how NSJF and OPT handle the jobs in $J_{\leq p}(t)$, though we measure progress differently for each:
  • For OPT, let $V_{\leq p}(t)$ denote the total workload of jobs in $J_{\leq p}(t)$ that has been processed by time $t$.
  • For NSJF, let $V^{\prime}_{\leq p}(t)$ denote the total processing time of jobs in $J_{\leq p}(t)$ that have started under NSJF by time $t$; that is, the total processing time of job $j$ in the set $\{j\in J_{\leq p}(t)\mid s_{j}(\textup{{NSJF}})\leq t\}$.
Lemma 27.
For every $t\geq 0$, define $t^{\prime}=A^{-1}(m(t+\tau))$. Then for every upper bound $p$, we have: $V^{\prime}_{\leq p}(t^{\prime})\geq V_{\leq p}(t).$
@@
Fix an arbitrary $p$, let $t_{0}$ be the last time some machine is idle by NSJF before $t^{\prime}$. Define $s_{j}$ as the latest start time, within the interval $(t_{0},\,t+\tau)$, at which NSJF schedules a job with $p_{j}>p$. If no such job exists, we set $s_{j}=t_{0}$ and conceptually assume that an infinitely large job starts at time $s_{j}$.
@@
We now analyze two cases:
Case 1: $s_{j}\geq t$. We straightforwardly have:
@@
Case 2: $s_{j}<t$. Then the time interval $s_{j}+\tau,t^{\prime})$ must be fully used by NSJF to process jobs with size at most $p$ (if any).
@@
In both cases, we conclude:
$V^{\prime}_{\leq p}(t^{\prime})\geq V_{\leq p}(t),$
which proves the lemma. ∎
Denote $N_{\leq p}(t)$ (or $N^{\prime}_{\leq p}(t)$ below) as the number of jobs in $J_{\leq p}(t)$ completed by time $t$ by OPT (or NSJF, respectively). Note that without the subscription $\leq p$, we remove the condition on job size. We prove:
Lemma 28.
For every $t\geq 0$, define $t^{\prime}=A^{-1}(m(t+2\tau))$. we have $N^{\prime}_{\leq p}(t^{\prime})\geq N_{\leq p}(t)$.
@@
Define $t_{1}=A^{-1}(m(t+\tau))$ and let $t_{2}=t^{\prime}$. By Lemma 27, we have the volume bound: $V^{\prime}_{\leq p}(t_{1})\geq V_{\leq p}(t)$ for every $p$.
We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t_{1}$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$. That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
To prove this, assume for contradiction that there exists the smallest index $i$ such that $p^{\prime}_{i}>p_{i}$. Under this assumption, we have the following bound for OPT:
$V_{\leq p_{i}}(t)\geq\sum_{j=1}^{i}p_{j},$
since at least $i$ jobs of size at most $p_{i}$ are completed by time $t$ in OPT.
Now consider NSJF. Since the $i$-th smallest started job under NSJF has a size strictly greater than $p_{i}$, all started jobs of size $\leq p_{i}$ must be among the first i1i-1 jobs. Therefore, we have:
@@
where the second inequality uses the assumption that $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t_{1})<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 27. Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
Finally, this implies that the number of started jobs with $p_{j}\leq p$ by time $t_{1}$ under NSJF is at least the number of completed jobs with $p_{j}\leq p$ by time $t$ under OPT. Moreover, since the number of active machines in NSJF is non-decreasing over time, all jobs started by time $t_{1}$ must be completed by time $t_{2}$. This completes the proof of the lemma.
@@
Let $u$ denote the maximum completion time under NSJF. The flow times of NSJF and OPT can be expressed as:
@@ Appendix B Randomized Non-Preemptive Algorithm: Multiple Machines
In this section, we extend Algorithm 2 from single-machine case to multi-machine case. The intuition is already introduced. We detail the dynamic but online stable algorithm in Algorithm 10.
@@
   Classify job $i$ large or small based on Partition($i,\ell$) with $\ell=\lfloor\sqrt{nm}\rfloor$;
 if $i$ $i$ is classified as large then
@@
There exists an online polynomial-time randomized non-preemptive algorithm for total flow time minimization, e.g., Algorithm 10, that is $O(\sqrt{n/m})$-competitive against the preemptive offline solution.
@@
The algorithm is online simulatable and runs in polynomial time, for the same reason as in Lemma 5. We use $S$ and $L$ to denote the sets of small and large jobs, where $S$ also includes proxy jobs. We use $J$ to denote the real job set, which includes only actual jobs and excludes proxy jobs.
@@
$F(\S_{1})=O\left(n\cdot\frac{1}{\sqrt{nm}}\right)\cdot F(\textup{{OPT}}(S))\leq O(\sqrt{n/m})\cdot\textup{{OPT}}.$
@@
$\mathbb{E}[|D_{k,l}|]=\sum_{k=1}^{\lfloor\sqrt{n/m}\rfloor}\sum_{l=1}^{m}|D_{k,l}|\cdot\Pr[w_{j}=k]\cdot\Pr[m_{j}=l]\leq O(\sqrt{n/m}),$
@@
For the self-delay, if $j$ is proxied by $j^{\prime}$, we count $r_{j^{\prime}}-r_{j}$; otherwise, we count $C^{\prime}_{j}-r_{j}$, where $C^{\prime}_{j}$ is the completion time of $j$ immediately after insertion. This self-delay is denoted by $\hat{F}_{j}$. Ignoring the final $p_{j}$ processing time, the delay term only comes from collecting idle time or passing through busy time.

### v3

@@ 8.3 Generalization to Multiple Machines with Kill-and-Restart
Any randomized online algorithm for total flow time minimization on parallel machines has a competitive ratio of at least $\Omega(\sqrt{n/m})$, for every fixed pair $(n,m)$, even with the capability of kill-and-restart.
@@
By Lemma 24, $\textup{{OPT}}=O(km)$, hence any deterministic algorithm is $\Omega(k)=\Omega(\sqrt{n/m})$ competitive and the theorem is proved. ∎
@@ 8.4 Deterministic Algorithms: Single Machine with Kill-and-Restart
    if ALG hasn’t killed job $u$ before time $t^{\prime}+\tau$ then
@@
Any deterministic online algorithm with kill-and-restart for total flow time minimization on a single machine has a competitive ratio of at least $\Omega(n/\log n)$.
@@ 8.5 Deterministic Algorithms: Multiple Machines without Kill-and-Restart
The lower bound $\Omega(n/m^{2}+\sqrt{n/m})$ follows by combining our randomized lower bound of $\Omega(\sqrt{n/m})$ with the deterministic $\Omega(n/m^{2})$ lower bound from [10], as presented below.
@@
Any deterministic online algorithm for total flow time minimization on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed pair $(n,m)$, where $\Gamma(n,m)=\Omega(n/m^{2}+\sqrt{n/m})$.
@@
We first show that for every $n$, $m$, we can construct an adversary to prove $\Omega(n/m^{2})$, the same as [10]. We still present it for completeness. Consider the following case for any deterministic algorithm ALG:
@@
OPT can start all small jobs at their release time immediately; it can also solve the first job at 0 if $t>1$, or solve the first job at $2$ otherwise. Hence, OPT is $O(1)$ in both cases. For ALG, as it can only use m1m-1 machines to solve the small jobs, at most m1m-1 jobs can be solved between the releases of two small job batches. The total flow time for ALG is at least $\sum_{i=1}^{n/m}i/n=\Omega(n/m^{2})$. Therefore, the competitive ratio is $\Omega(n/m^{2})$.
@@ 9 Lower Bounds for Algorithms without the Knowledge of $n$
Adversary construction. We construct an oblivious adversary as follows. At time 0, the adversary releases a single job of size $2$. Let $p(t)$ denote the probability density that ALG starts this job at time $t$, assuming no other jobs are present. Since $p(t)$ depends only on the algorithm’s description, it is known to the adversary in advance. The algorithm must fall into one of the following categories (the parameters $x$ and $y$ will be fixed later):
@@
    For all $t\in\{1,2,\dots,x\}$, it holds that $\int_{t}^{t+1}p(s)\,ds<y$. In this case, the probability that the size-$2$ job has not been started by time $x$ is at least 1xy1-xy. The adversary releases $n_{0}$ dummy jobs of size 0 at time $x+1$. If ALG delays the size-$2$ job until after time $x$, its flow time alone is at least $x+2$, while $\textup{{OPT}}\leq 2$. Therefore,
@@
The $\sqrt{n/m}$ term follows from the known-$n$ lower bound (Theorem 5). For the $n/m^{2}$ term, the intuition resembles Lemma 9: losing one machine out of $m$ incurs an $\Omega(n/m^{2})$ penalty in flow time.
@@
Since the $\Omega(\sqrt{n/m})$ lower bound holds even when $n$ is known (Theorem 5), it suffices to construct a hard instance achieving ratio $\Omega(n/m^{2})$.
Adversary construction. We use the same setup as in the single-machine case: the adversary first releases a size-2 job at time 0 and classifies the algorithm into Type A or Type B based on the density $p(t)$. The only difference lies in how we construct the hard instance for Type A.
  1. A.
    If ALG is of Type A, the adversary releases $n_{1}$ additional jobs of size $1/k$, where $k=n_{1}/m$. The release pattern is:
@@
Appendix A Generalized Version of Non-Preemptive Shortest Job First
@@
For any processing time threshold $p$ and time $t$, let $J_{\leq p}(t)$ be the set of jobs with $p_{j}\leq p$ and release time $r_{j}\leq t$.
Step 1: Volume Matching.
For each algorithm, we track the volume of work associated with jobs in $J_{\leq p}(t)$. Let $V_{\leq p}(t)$ denote the total volume of work processed on jobs in $J_{\leq p}(t)$ by time $t$ under OPT, and let $V^{\prime}_{\leq p}(t)$ denote the total volume of jobs in $J_{\leq p}(t)$ that have been started by time $t$ under NSJF.
Lemma 27 (Volume Matching).
For every $t\geq 0$, define $t^{\prime}=A^{-1}(m(t+\tau))$. Then for every threshold $p$, we have $V^{\prime}_{\leq p}(t^{\prime})\geq V_{\leq p}(t)$.
@@
Fix an arbitrary $p$, and let $t_{0}$ be the last time some machine is idle under NSJF before $t^{\prime}$. Let $j$ be the job with $p_{j}>p$ that is started latest by NSJF within the interval $(t_{0},t^{\prime})$. If no such job exists, we let $s_{j}=t_{0}$ for a fictitious job $j$ of infinite size.
@@
We now analyze two cases. If $s_{j}\geq t$, we immediately have:
@@
Otherwise, if $s_{j}<t$, then since $s_{j}$ is the latest start of a job with $p_{j}>p$ after the last idle time $t_{0}$, all processing in $s_{j}+\tau,t^{\prime})$ is on jobs of size at most $p$, and no machine is idle in this interval. Therefore:
@@
Step 2: Completion Matching.
Let $N_{\leq p}(t)$ and $N^{\prime}_{\leq p}(t)$ denote the number of jobs in $J_{\leq p}(t)$ completed by time $t$ under OPT and NSJF, respectively. When the subscript $\leq p$} is omitted, we count all completed jobs regardless of size.
Lemma 28 (Completion Matching).
For every $t\geq 0$, define $t^{\prime}=A^{-1}(m(t+2\tau))$. For every threshold $p$, we have $N^{\prime}_{\leq p}(t^{\prime})\geq N_{\leq p}(t)$.
@@
Define $t_{1}=A^{-1}(m(t+\tau))$.
For each rank $i$, let $p_{i}$ be the size of the $i$-th smallest job completed under OPT by time $t$, and let $p^{\prime}_{i}$ be the size of the $i$-th smallest job started under NSJF by time $t_{1}$. Let $k$ and $k^{\prime}$ denote the total number of such jobs, respectively. We claim that $k^{\prime}\geq k$ and $p^{\prime}_{i}\leq p_{i}$ for all $i\leq k$.
Assume for contradiction that $i\leq k$ is the smallest rank such that either $i>k^{\prime}$ or $p^{\prime}_{i}>p_{i}$. For OPT, at least $i$ jobs of size at most $p_{i}$ are completed by time $t$, so:
$V_{\leq p_{i}}(t)\geq\sum_{j=1}^{i}p_{j}.$
For NSJF, all started jobs of size $\leq p_{i}$ are among the first i1i-1 jobs (either because only i1i-1 jobs have started, or because the $i$-th smallest is strictly larger than $p_{i}$). Therefore:
@@
where the second inequality uses $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. This gives $V^{\prime}_{\leq p_{i}}(t_{1})<V_{\leq p_{i}}(t)$, contradicting Lemma 27.
Therefore $k^{\prime}\geq k$ and $p^{\prime}_{i}\leq p_{i}$ for all $i\leq k$, which implies that for every threshold $p$, the number of jobs of size $\leq p$ started by NSJF by time $t_{1}$ is at least the number completed by OPT by time $t$. Moreover, since the number of active machines in NSJF is non-decreasing over time, all jobs started by time $t_{1}$ are completed by time $t^{\prime}$. This completes the proof. ∎
Step 3: Bounding the Flow Time.
@@
Let $u$ denote the maximum completion time under NSJF. The total flow times can be expressed as:
@@ Appendix B Randomized Non-Preemptive Algorithm: Multiple Machines
In this section, we extend Algorithm 2 from the single-machine case to the multi-machine case. The intuition is introduced in Section 4. We detail the dynamic but online stable algorithm in Algorithm 10.
@@
   Call Partition($i,\ell$) with $\ell=\lfloor\sqrt{nm}\rfloor$ to classify job $i$ and update related jobs;
 if $i$ is classified as large then
@@
Algorithm 10 is an online polynomial-time randomized non-preemptive algorithm for total flow time minimization that is $O(\sqrt{n/m})$-competitive against the preemptive offline solution.
@@
The algorithm is online stable and runs in polynomial time, for the same reason as in Lemma 5. We use $S$ and $L$ to denote the sets of small and large jobs, where $S$ also includes proxy jobs. We use $J$ to denote the real job set, which includes only actual jobs and excludes proxy jobs.
@@
$F(\S_{1})=O(\sqrt{n/m})\cdot F(\textup{{OPT}}(S))=O(\sqrt{n/m})\cdot\textup{{OPT}}.$
@@
$\mathbb{E}[|D_{w_{j},m_{j}}|]=\sum_{k=1}^{\lfloor\sqrt{n/m}\rfloor}\sum_{l=1}^{m}|D_{k,l}|\cdot\Pr[w_{j}=k]\cdot\Pr[m_{j}=l]=O(\sqrt{n/m}),$
@@
For the self-delay, if $j$ is proxied by $j^{\prime}$, we count $r_{j^{\prime}}-r_{j}$; otherwise, we count $\hat{C}_{j}-r_{j}$, where $\hat{C}_{j}$ is the completion time of $j$ immediately after insertion. This self-delay is denoted by $\hat{F}_{j}$. Ignoring the final $p_{j}$ processing time, the delay term only comes from collecting idle time or passing through busy time.